There are currently 73 million cars in a certaint country increasing by 1.7% annually how many years will it take for this country to reach 91 million cars? Round to the nearest year
Answer by FrankM(1040)

73 * 1.017^N = 91 < note, an increase of 1.7% is the same as multiplying by 1.017

1.017^N = 91/73 or 1.2466

log of both sides
log 1.017^N = log 1.2466
rearrange to
N log 1.017 = log 1.2466
N= log 1.2466/log 1.017
N= 13.075
Ask your teacher if you round down to nearest integer, or up to 14, the year sales actually exceeded 91 M